OpenStudy (alyssajobug):

What is the solution to: √2x+13 -5=x? x = -13 x = -10/11 x = 6 x = 3/7

OpenStudy (anonymous):

√2x+13 -5=x we first want to isolate all the x terms to one side. so can you move the x's to the same side and move all the 'non x's' to the other side using algebra?

OpenStudy (anonymous):

Is that \[\sqrt{2x+13}-5=x~~?\]

OpenStudy (alyssajobug):

yes

OpenStudy (alyssajobug):

@SithsAndGiggles Yes it is

OpenStudy (anonymous):

i'd move the 5 over to the other side, then square both sides to get rid of the square root sqrt(2x+13) - 5 = x sqrt(2x+13) = x + 5 (sqrt(2x+13))^2 = (x+5)^2 2x + 13 = (x+5)^2